In 2006, the number of federal hazardous waste sites in California was 5 less than twice the number of sites in Washington. How many hazardous waste sites were there in Washington if there were 27 such sites in California?
Answer by mengayzkie(3)

Given:
Let x = the # of hazardous sites in California
y = the # of hazardous sites in Washington
The number of hazardous waste sites in California can be represented as:
x = 2y - 5 which is equal to 27
Substitute the value of x from the given, thus giving us:
27 = 2y - 5
27 + 5 = 2y
to solve for y:
32= 2y
y = 16
therefore there are 16 hazardous waste sites in Washington
checking:
substitute the values of x and y:
27 = 2(16) - 5
27 = 32 - 5
27 = 27 checked!